Let ?: ?2(R) ⟶ R3 be a linear transformation such that ?(1) = (−1, 2, −3), ?(1 + 3?) = (4, −5, 6), and ?(1 + ?2) = (−7, 8, −9).
a. Show that {1,1 + 3? ,1 + ?2} is a basis for ??2(R) (7pts)
b. Compute ?(−1 + 4? + 2?2). (3pts)
Answer #1

we we know that {1,2,x?} is a basis of Now } [(+sx) - ] - = x 2 = ç (173X) + 6).1 хе 之 span 1,1+3, Ita? also Ita? x (1+x²) +So 2 -1 +420 +22= -13 + (1+32) + 2 (1+x2) So T(-1+4% +224) = -13 TOWT(1+34) + 2 T T (2+x²) = - 13 (-1,2,-3) + 4/5 (9,-516) +